OpenStudy (anonymous):

5/8 x + 1/16 x =9/16+x the solution is x=???

OpenStudy (anonymous):

16[5/8x+1/16x=9/16+x]16 10x+x=9+16x 11x-16x=9 -5x=9 x=-9/5

OpenStudy (anonymous):

thanks :)